# Uncertainty Relation for Entropy and Temperature of Gibbs States

### 17 Mar 2026 v1

@@ II Entropy Fisher information
This bound is saturated asymptotically by a projective energy measurement (see Appendix B).
Equation (2) indicates a duality: the heat capacity governs the precision of $S$- and $T$-estimation inversely. Having a large $C_{\!v}$ allows for precise temperature estimation ($F_{T}\propto C_{\!v}$) as small temperature differences produce large, distinguishable shifts in the energy distribution. Conversely, large $C_{\!v}$ makes entropy hard to estimate ($F_{S}\propto 1/C_{\!v}$) because many distinct entropy values map to nearly identical energy statistics.
@@ III Universal uncertainty relation
Crucially, this bound is saturatable: because projective energy measurement simultaneously saturates both individual Cramér–Rao bounds, the product $\Delta^{2}S\,\Delta^{2}T=T^{2}/n^{2}$ is achievable, distinguishing Eq. (5) from a generic (and potentially vacuous) product of independent lower bounds. Equation (5) is the metrological signature of thermodynamic conjugacy: for any pair of Legendre-conjugate variables, an analogous relation holds in which all response functions cancel (see Sec. VIII).
@@ IV Two-level system
Figure 1: Fisher information for entropy and temperature estimation in (a) a two-level system with energy gap $\Delta$ and (b) a quantum harmonic oscillator with frequency $\omega$. Solid curves show $F_{S}=1/C_{\!v}$ (blue) and $F_{T}=C_{\!v}/T^{2}$ (red); the dashed curve is their product $F_{S}\cdot F_{T}=1/T^{2}$. In panel (b), green dots mark the two crossovers at $C_{\!v}=T$ ($T\approx 0.22\,\hbar\omega$ and $T\approx 0.90\,\hbar\omega$); the shaded region indicates the intermediate “thawing” regime where $F_{T}>F_{S}$. The dotted line marks the classical equipartition floor $F_{S}\to 1$.
@@ VII Relation to thermodynamic geometry
Ruppeiner [18, 19] defined a Riemannian metric on equilibrium states via $g^{R}_{UU}=-\partial^{2}S/\partial U^{2}=1/(T^{2}C_{\!v})$. Janyszek and Mrugała [9] showed that this coincides with the Fisher information metric for Gibbs distributions. Under the coordinate change dU=TdSdU=T\,dS, the metric transforms to $g^{R}_{SS}=1/C_{\!v}=F_{S}$. Thus $F_{S}$ is precisely the Ruppeiner metric component in entropy coordinates—an identification that is implicit in the information-geometry literature but whose estimation-theoretic content has not previously been exploited.
@@ VIII Other conjugate pairs
The argument of Sec. II extends to every Legendre-conjugate pair. In each case the relevant Gibbs state is an exponential family in the intensive parameter, the QFI for that parameter equals $\beta^{2}$ times the variance of its conjugate operator, and the Fisher information for the extensive variable follows by reparametrisation. The conjugate susceptibility cancels in the product.
@@
with the isothermal compressibility $\kappa_{T}$ cancelling identically from the product (see Appendix F for the full derivation). The Cramér–Rao bound gives $\Delta^{2}V\;\Delta^{2}P\geq T^{2}/n^{2}$.
@@ IX Rényi entropy Fisher information
The von Neumann entropy $S=S_{1}$ is distinguished among the Rényi family $S_{\alpha}=(1{-}\alpha)^{-1}\ln\mathrm{Tr}(\rho^{\alpha})$ by its direct connection to thermodynamic conjugacy. For a Gibbs state, one may show (Appendix G) that $S_{\alpha}(\beta)$ is strictly monotone for every $\alpha>0$, so the reparametrisation argument applies, yielding the Rényi entropy Fisher information
@@ X Generalisations
Grand canonical ensemble.—For $\rho(\beta,\mu)=e^{-\beta(H-\mu N)}/\mathcal{Z}$, projecting the two-parameter Fisher matrix onto the entropy direction (see Appendix C) yields
@@
Generalised Gibbs ensemble.—For $\rho=e^{-\sum_{k}\lambda_{k}I_{k}}/Z$ with mutually commuting conserved charges $\{I_{k}\}$ and conjugate Lagrange multipliers $\{\lambda_{k}\}$, the entropy $S=\sum_{k}\lambda_{k}\langle I_{k}\rangle+\ln Z$ can be estimated by projecting the multiparameter Fisher matrix $F]_{kl}=\mathrm{Cov}(I_{k},I_{l})$ onto the entropy gradient (see Appendix D), giving
@@ Acknowledgements.
Acknowledgments We would like to thank George Mihailescu and Paulo Abiuso for their comments and feedback on the draft. We would like to thank Shubhang Dadhich, Kjartan van Driel, Ned O’Reilly, Marc Nairn, Mathias Bo Svendsen, Parvinder Solanki and particularly Enrique Verduras for fruitful discussions.
@@ Appendix G Rényi entropy Fisher information: detailed derivation
The Rényi entropy Fisher information follows by the same covariant transformation used in Sec. II, yielding Eq. (14).

### 18 Mar 2026 v2

@@ II Entropy Fisher information
This bound is saturated asymptotically by a projective energy measurement (see Appendix). Equation (2) indicates a duality: the heat capacity governs the precision of $S$- and $T$-estimation inversely. Having a large $C_{\!v}$ allows for precise temperature estimation ($F_{T}\propto C_{\!v}$) as small temperature differences produce large, distinguishable shifts in the energy distribution. Conversely, large $C_{\!v}$ makes entropy hard to estimate ($F_{S}\propto 1/C_{\!v}$) because many distinct entropy values map to nearly identical energy statistics.
@@ III Universal uncertainty relation
Crucially, this bound is saturatable: because projective energy measurement simultaneously saturates both individual Cramér–Rao bounds, the product $\Delta^{2}S\,\Delta^{2}T=T^{2}/n^{2}$ is achievable, distinguishing Eq. (5) from a generic (and potentially vacuous) product of independent lower bounds. Equation (5) is the metrological signature of thermodynamic conjugacy: for any pair of Legendre-conjugate variables, an analogous relation holds in which all response functions cancel.
@@ IV Two-level system
Figure 1: Fisher information for entropy and temperature estimation in (a) a two-level system with energy gap $\Delta$ and (b) a quantum harmonic oscillator with frequency $\omega$. Solid curves show $F_{S}=1/C_{\!v}$ (blue) and $F_{T}=C_{\!v}/T^{2}$ (red); the dashed curve is their product $F_{S}\cdot F_{T}=1/T^{2}$.
@@ VII Relation to thermodynamic geometry
Ruppeiner [18, 19] defined a Riemannian metric on equilibrium states via $g^{R}_{UU}=-\partial^{2}S/\partial U^{2}=1/(T^{2}C_{\!v})$. Janyszek and Mrugała [9] showed that this coincides with the Fisher information metric for Gibbs distributions. Under the coordinate change dU=TdSdU=T\,dS, the metric transforms to $g^{R}_{SS}=1/C_{\!v}=F_{S}$. Thus $F_{S}$ is precisely the Ruppeiner metric component in entropy coordinates, an identification that is implicit in the information-geometry literature but whose estimation-theoretic content has not previously been exploited.
@@ VIII Other conjugate pairs
Our main argument extends to every Legendre-conjugate pair. In each case the relevant Gibbs state is an exponential family in the intensive parameter, the QFI for that parameter equals $\beta^{2}$ times the variance of its conjugate operator, and the Fisher information for the extensive variable follows by reparametrisation. The conjugate susceptibility cancels in the product.
@@
with the isothermal compressibility $\kappa_{T}$ cancelling identically from the product (see Appendix for the full derivation). The Cramér–Rao bound gives $\Delta^{2}V\;\Delta^{2}P\geq T^{2}/n^{2}$.
@@ IX Rényi entropy Fisher information
The von Neumann entropy $S=S_{1}$ is distinguished among the Rényi family $S_{\alpha}=(1{-}\alpha)^{-1}\ln\mathrm{Tr}(\rho^{\alpha})$ by its direct connection to thermodynamic conjugacy. For a Gibbs state, one may show (see Appendix) that $S_{\alpha}(\beta)$ is strictly monotone for every $\alpha>0$, so the reparametrisation argument applies, yielding the Rényi entropy Fisher information
@@ X Generalisations
Grand canonical ensemble.—For $\rho(\beta,\mu)=e^{-\beta(H-\mu N)}/\mathcal{Z}$, projecting the two-parameter Fisher matrix onto the entropy direction (see Appendix) yields
@@
Generalised Gibbs ensemble.—For $\rho=e^{-\sum_{k}\lambda_{k}I_{k}}/Z$ with mutually commuting conserved charges $\{I_{k}\}$ and conjugate Lagrange multipliers $\{\lambda_{k}\}$, the entropy $S=\sum_{k}\lambda_{k}\langle I_{k}\rangle+\ln Z$ can be estimated by projecting the multiparameter Fisher matrix $F]_{kl}=\mathrm{Cov}(I_{k},I_{l})$ onto the entropy gradient (see Appendix), giving
@@ Acknowledgements.
Acknowledgments We would like to thank George Mihailescu, Paolo Abiuso and Daniel Braun for their comments and feedback on the draft. We would like to thank Shubhang Dadhich, Kjartan van Driel, Ned O’Reilly, Marc Nairn, Mathias Bo Svendsen, Parvinder Solanki and particularly Enrique Verduras for fruitful discussions.
@@ Appendix G Rényi entropy Fisher information: detailed derivation
The Rényi entropy Fisher information follows by the same covariant transformation used in the , yielding Eq. (14).